OpenStudy (anonymous):

solve for x with the 3 equations in a triangle 6x+2 2x+5 x+11 no crazy stuff, simply explain.

OpenStudy (turingtest):

6x+2=2x+5=x+11 ?????????????????? is that what you mean by triangle?